OpenStudy (anonymous):

Find the domain and range of f(x0 = (x^2-9)/(x-3)

OpenStudy (anonymous):

Domain = R-{3}

OpenStudy (anonymous):

\[f(x)=\frac{(x-3)(x+3)}{(x-3)}\] and x can not be 3 why?

OpenStudy (anonymous):

then range

OpenStudy (anonymous):

ok range f(x)=x+3 am i right?

OpenStudy (anonymous):

so range is all real numbers??

OpenStudy (anonymous):

no because u cant let x=3

OpenStudy (anonymous):

Range = R-{6}

OpenStudy (anonymous):

hw

OpenStudy (anonymous):

f(3)=3+3=6

OpenStudy (anonymous):



OpenStudy (anonymous):

i got it thxxx

OpenStudy (anonymous):

plzz take my new question